Question 1.13 [2, 2, 3]


Due to COVID-19 there was no time to have the swimming gala at your old primary school. The


principal knows that you are currently studying statistics and he wants you to help with this


probability problem. The principal tells you that out of the 8 swimmers, 3 are from grade 4, 2 are


from grade 5 and 3 are from grade 6. Since no Gala can be held the principal selects swimmers at


random to attend the EP school Gala, the first student selected at random will be representing the


schools fastest swimmer, while the second student selected will represent the school second fastest


swimmer.


Help the principal to answer the following questions:


a) What is the probability that the two fastest swimmers are from Grade 6?


b) What is the probability that fastest swimmer is from Grade 4 and the second fastest from


Grade 6?


c) Timothy is a student in grade 5, what is the probability that he will either come first or second?

1
Expert's answer
2022-06-13T10:02:06-0400

a)


"P(6\&6)=\dfrac{3}{8}(\dfrac{2}{7})=\dfrac{3}{28}"

b)


"P(4\&6)=\dfrac{3}{8}(\dfrac{3}{7})+\dfrac{3}{8}(\dfrac{3}{7})=\dfrac{9}{28}"

c)


"P(Timothy)=\dfrac{\dbinom{7}{1}}{\dbinom{8}{2}}=\dfrac{1}{4}"